# Towards Quantifying Long-Range Interactions in Graph Machine Learning: a Large Graph Dataset and a Measurement

### v1

@@ 6 Theoretical Justification of Jacobian-Based Measurement
Lastly, in this section, we provide the intuition behind the proposed measurement for quantifying long-rangeness in Section 4. In particular, we discuss potential limiting cases and provide a theoretical understanding of influence score dilution for lattice and grid-like graphs, which supports the construction of the proposed measurement. The proofs for all lemmata, propositions, and theorems can be found in Appendix D.2.
6.1 Intuition behind Jacobian-based Influence Score
where $\bm{X}$ is the original (unperturbed) input feature matrix, and $\delta\bm{e}_{uj}$ is an infinitesimal perturbation in the $j^{th}$ component of the feature vector at node $u$ (a standard basis vector in the feature space of the node). This means the more positive the Jacobian entry is, the more a positive perturbation of the features at node $u$ and component $j$ will increase the logits at node $v$ and component $i$ at the final layer. In other words, the Jacobian entry being positive or negative means that the logits are pushed up or down. Given that, after applying the softmax function, the probabilities at a given data point increase monotonically with the logits at that point [^4] It is well-known that for a vector of logits the softmax probabilities are monotonically increasing with respect to their corresponding logits. Also see Appendix D.2 for a proof., we can consider both positive and negative influences as actual influence and only need to focus on the absolute value (i.e., sensitivity rather than direction).
6.2 Potential Limitations: Output Cancellation
Proposition 6.1 (Absolute Jacobian sensitivity may over-estimate influence).
There exists a model $k$ with zero net derivative (i.e., $k^{\prime}(\bm{X})=0\,~{}\forall\bm{X}$) for which the sum of the absolute values of individual Jacobian contributions is nonzero.
The proof relies on constructing a toy function with two branches that counteract each other exactly for any input, see Appendix D.2. The nonzero value in the proposition suggests a strong sensitivity to perturbations in $\bm{X}$ even though the net change in the output $k(\bm{X})$ is zero. Thus, this counterexample demonstrates that measuring influence solely via the absolute values of the Jacobian entries can be misleading when opposing contributions cancel in the final output. Although this is a valid concern, we next show that for Message Passing Neural Networks (MPNNs), at least at initialization, such cancellation does not happen.
Definition 6.2 (Message-Passing Neural Network layer).
Theorem 6.3 (Measure-zero of exact cancellation at MPNN initialization).
6.3 Justification of Proposed Measurement and the Mean Influence Score Dilution Problem
Finally, we justify the proposed measurement for analyzing long-range interactions in GNNs, particularly by comparing the mean and the sum of influence scores of neighboring nodes as candidates to represent the total influence score for a focal node $v$, $T_{h}(v)$. Recall that $T_{h}(v)=I_{\text{sum}}(v,h)=\sum_{u:\rho(v,u)=h}I(v,u)$ as in Equation (3). We first revise important notions such as that of the standard lattice in Euclidean space, $h$-hop shells, and quasi-isometric graphs, and then provide a mathematical analysis of influence score dilution. This helps justify the construction of $T_{h}(v)$ as a sum rather than a mean operator, which is key to the computation of the average total influence in Equation (3) and the influence-weighted receptive field in Equation (4).
Definition 6.4 (Standard lattice in $\mathbb{R}^{D}$).
Definition 6.5 ($h$-hop shells).
Definition 6.6 (Quasi-isometric graphs).
@@ 6.3 Justification of Proposed Measurement and the Mean Influence Score Dilution Problem
Given the definitions above, and after having quantified the growth of the $h$-hop shell, one can prove the following theorem, which motivates our aggregation choice.
@@
Suppose that for a fixed $v$ and for each $h\geq h_{0}$ there exists a distinguished node $u^{*}\in\mathcal{N}_{h}(v)$ with a strong influence on $v$, quantified by $I(v,u^{*})=I^{*}>0,$ while for all other nodes $u\in\mathcal{N}_{h}(v)\setminus\{u^{*}\}$ the influence $I(v,u)$ is negligible (or zero). Define $I_{\text{sum}}(v,h)=\sum_{u\in\mathcal{N}_{h}(v)}I(v,u)=\sum_{u:\rho(v,u)=h}I(v,u)\quad\text{and}\quad I_{\text{mean}}(v,h)=\frac{1}{|\mathcal{N}_{h}(v)|}\sum_{u\in\mathcal{N}_{h}(v)}I(v,u).$ Then, $I_{\text{sum}}(v,h)\geq I^{*}$, and $I_{\text{mean}}(v,h)\leq\dfrac{I^{*}}{\mathfrak{C}_{1}\,h^{D-1}}$. Hence, as $h\to\infty$, we have $I_{\text{mean}}(v,h)\to 0,$ while $I_{\text{sum}}(v,h)$ remains bounded below by $I^{*}$. This also holds for planar graphs, where $D=2$.
@@
Let $\mathcal{G}$ be a grid-like graph in $D$ dimensions, and define the aggregated $h$-hop neighborhood (or ball) of a node $v$ as $B_{h}(v)=\bigcup_{i=1}^{h}\mathcal{N}_{i}(v).$ As before, suppose that within $B_{h}(v)$ there exists a unique node $u^{*}$ with influence $I(u^{*},v)=I^{*}>0$ and that for all other nodes $u\in B_{h}(v)\setminus\{u^{*}\}$, the influence is negligible. Then, the mean aggregated influence is diluted at a rate proportional to $1/h^{D}$. In particular, for a planar graph ($D=2$), the dilution occurs at a rate proportional to $1/h^{2}$.
Hence, this analysis formally justifies the choice in Equation (3), which considers the aggregate influence of neighboring nodes $T_{h}(v)=I_{\text{sum}}$ as a more reliable measure than the mean $I_{\text{mean}}$, which is susceptible to dilution, particularly in the case of distant neighbors. Finally, note the following:
@@ 7 Conclusion
The main objective of our work is to provide better tools to help quantify long-range interactions in GNNs. Previous benchmarks, such as the LRGB (Dwivedi et al., 2023), are introduced in the context of small graph inductive learning, using solely the performance gap between classical GNNs and Graph Transformers to support the presence of long-range signals. In this work, we introduce a new large graph dataset based on city road networks, featuring long-range dependencies for transductive learning, and propose a principled measurement to quantify such dependencies. We also provide theoretical justification of both the proposed dataset and measurement, focusing on over-smoothing and influence score dilution. Beyond benchmarking purposes, our work also holds potential for a broader impact, informing applications in urban planning and transportation by providing tools to analyze and understand accessibility and influence within city road networks.
Appendix A Dataset Details
A.1 Network Statistics
@@
Since the exact calculation of $D$ has a computational complexity $\mathcal{O}(N^{2}\log(N))$, we use the following approach to estimate the approximate diameter $\hat{D}$ of our city networks. For all nodes in a given city, we select the ones with the maximum and minimum latitude and longitude, respectively: $coord(v_{1})=(\;\cdot\;,\;\text{lat}_{max})$, $coord(v_{2})=(\;\cdot\;,\;\text{lat}_{min})$, $coord(v_{3})=(\text{long}_{max},\;\cdot\;)$, and $coord(v_{4})=(\text{long}_{min},\;\cdot\;)$. Based on these, we compute the shortest path between $(v_{1},v_{2})$ and $(v_{3},v_{4})$, and take their maximum as our final diameter estimate, that is, $\hat{D}=\max\big{(}d(v_{1},v_{2}),d(v_{3},v_{4})\big{)}$. Note that the exact diameter $D$ will always be larger than our estimation $\hat{D}$.
@@
A.2 Feature and Label Construction
@@
Based on the raw city map queries, we obtained the following features on junctions and road segments:
@@
Since the categorical features land use, lanes, and road type contain many categories of only a few data points and varies across different networks, we only take the top $8$ categories with most entries for each categorical feature, and treat the rest as a single category - other. Based on our observations, we can cover more than 90% of the network with the top $8$ categories in all cases. This strategy leads to $12$ node features and $25$ edge features after one-hot encoding, that is, a total of $37$ features after transferring edge features to their adjacent nodes by neighborhood aggregation. Finally, we transform the graph into an undirected one by merging all edges between each pair of nodes into a single edge, where the edge features are averaged using to_undirected(reduce="mean") from PyG.
Node labels based on eccentricity estimation.
As mentioned in Section 2, we use a $16$-hop ego-network to estimate the eccentricity of each node. Specifically, after obtaining the neighborhood $\mathcal{N}_{16}(v)$ within 16 hops of node $v$, we compute the shortest path from node $v$ to all nodes within this neighborhood, and take the maximum as the estimated eccentricity $\hat{\varepsilon}(v)$:
$\hat{\varepsilon}(v)=\max_{u\in\mathcal{N}_{16}(v)}\rho_{w}(v,u),\quad\rho_{w}(v,u)=\min_{\pi\in P(v,u)}\sum_{e\in\pi}w(e),$ (24)
@@
where $\rho_{w}(v,u)$ is the weighted shortest path distance from node $v$ to node $u$, $P(v,u)$ denotes the set of all possible paths from node $v$ to node $u$, and $w(e)$ represents the edge weight for edge $e$. Here, we use one of the edge features road length as the edge weight $w(e)$, such that the estimated eccentricity $\hat{\varepsilon}(v)$ will indicate the maximal traveling distance (in meters) from node $v$ to its $16$-th hop neighbors. Figure 6 shows the distributions of such approximation across different city networks, in which Paris and Shanghai have the most skewed and uniform distributions, respectively. After obtaining the approximations for all nodes, we split them into $10$ quantiles as the final node labels.
Then, the computation is at a much cheaper cost of $\mathcal{O}(|\hat{E}|+|\hat{V}|\log|\hat{V}|)$ using Dijkstra’s algorithm (Dijkstra, 1959), where $|\hat{V}|$ and $|\hat{E}|$ are the average number of nodes and edges across all 16-hop ego-networks. The calculation is implemented with networkx (Hagberg et al., 2008) on a CPU cluster of $72$ Intel Xeon @ 2.3GHz cores, which takes 3, 5, 6, and 23 hours to compute for Paris, Shanghai, L.A. and London, respectively.
Appendix B Experiment Details
Figure 7: The parameter size of baselines under different number of layers $L\in[2,4,8,16]$.
B.1 Baseline Configurations and Resource Consumptions
Experimental setups.
As discussed in Section 3, we consider two scenarios to show the long-range dependency in our City-Networks:
  • Setting the number of layers $L$ equal to the number of sampling hops $H$, and then consider $L=H\in\{2,4,8,16\}$.
  • Fixing the number of layers at $L=16$, and then vary the number of sampling hops at $H\in\{2,4,8,16\}$.
For fair comparison purposes, we set the hidden channel dimension to $32$ for all models. Since the number of node features (37) and number of classes (10) are identical across networks, the parameter sizes of the model, as presented in Figure 7, will also remain unchanged on different networks. Note that although some models generally have more parameters than others, with the same number of layers, the parameter sizes are relatively at the same order across all models.
Table 5: Summary of hyper-parameters used for training.
train/valid/test # epochs lr weight decay dropout batch size record window hidden size
10%/10%/80% $20k$ $10^{-3}$ $10^{-5}$ $0.2$ $20k$ $100$ $32$
The hyperparameters used for training are summarized in Table 5. In particular, we run each model for $20k$ epochs with a learning rate of $10^{-3}$ for sufficient training, during which we record the validation accuracy every $100$ epochs and save the model at the best validation epoch for final testing. All the scenarios are repeated 5 times and we report their means and standard deviations.

### v2

@@ 6 Theoretical Justification of Jacobian-Based Measurement
Finally, we justify the proposed measurement for analyzing long-range interactions in GNNs, particularly by comparing the mean and the sum of influence scores of neighboring nodes as candidates to represent the total influence score for a focal node $v$, $T_{h}(v)$. Recall that $T_{h}(v)=I_{\text{sum}}(v,h)=\sum_{u:\rho(v,u)=h}I(v,u)$ as in Equation (2). We first introduce definitions for the standard lattice in Euclidean space, $h$-hop shells, and quasi-isometric graphs, and then provide a mathematical analysis of influence score dilution. Due to space limits, we provide the definitions in Appendix H.
@@
After having quantified the growth of the $h$-hop shell, one can prove the following theorem, which motivates our aggregation choice.
@@
Suppose that for a fixed $v$ and for each $h\geq h_{0}$ there exists a distinguished node $u^{*}\in\mathcal{N}_{h}(v)$ with a strong influence on $v$, quantified by $I(v,u^{*})=I^{*}>0,$ while for all other nodes $u\in\mathcal{N}_{h}(v)\setminus\{u^{*}\}$ the influence $I(v,u)$ is negligible (i.e. zero). Define $I_{\text{sum}}(v,h)=\sum_{u\in\mathcal{N}_{h}(v)}I(v,u)=\sum_{u:\rho(v,u)=h}I(v,u)\quad\text{and}\quad I_{\text{mean}}(v,h)=\frac{1}{|\mathcal{N}_{h}(v)|}\sum_{u\in\mathcal{N}_{h}(v)}I(v,u).$ Then, $I_{\text{sum}}(v,h)\geq I^{*}$, and $I_{\text{mean}}(v,h)\leq\dfrac{I^{*}}{\mathfrak{C}_{1}\,h^{D-1}}$. Hence, as $h\to\infty$, we have $I_{\text{mean}}(v,h)\to 0,$ while $I_{\text{sum}}(v,h)$ remains bounded below by $I^{*}$. This also holds for planar graphs, where $D=2$.
@@
Let $\mathcal{G}$ be a grid-like graph in $D$ dimensions, and define the aggregated $h$-hop neighborhood (or ball) of a node $v$ as $B_{h}(v)=\bigcup_{i=1}^{h}\mathcal{N}_{i}(v).$ As before, suppose that within $B_{h}(v)$ there exists a unique node $u^{*}$ with influence $I(u^{*},v)=I^{*}>0$ and that for all other nodes $u\in B_{h}(v)\setminus\{u^{*}\}$, the influence is negligible (i.e. zero). Then, the mean aggregated influence is diluted at a rate proportional to $1/h^{D}$. In particular, for a planar graph ($D=2$), the dilution occurs at a rate proportional to $1/h^{2}$.
The analysis above formally justifies the choice in Equation (2), which considers the aggregate influence of neighboring nodes $T_{h}(v)=I_{\text{sum}}$ as a more reliable measure than the mean $I_{\text{mean}}$, which is susceptible to dilution, particularly in the case of distant neighbors. This is key to the computation of the average total influence in Equation (2) and the influence-weighted receptive field in Equation (3). Finally, note the following:
@@
Proofs are provided in Appendix I, and a more detailed justification of Jacobian-based influence score and connections to existing literature, along with its limitations, can be found in Appendix G.
@@ 7 Conclusion
The main objective of our work is to provide better tools to help quantify long-range interactions in GNNs. Previous benchmarks, such as the LRGB (Dwivedi et al., 2023), are introduced in the context of small graph inductive learning, using solely the performance gap between classical GNNs and GTs to support the presence of long-range signals. In this work, we introduce a new large graph dataset based on city road networks, featuring long-range dependencies for transductive learning, and propose a principled measurement to quantify such dependencies. We also provide theoretical justification of both the proposed dataset and measurement, focusing on over-smoothing and influence score dilution. Beyond benchmarking purposes, our work also holds potential for a broader impact, informing applications in urban planning and transportation by providing tools to analyze and predict accessibility within city road networks.
Dataset and Code Availability
Our dataset, CityNetwork, is now publicly available in the latest version of PyTorch Geometric (PyG) as an InMemoryDataset under torch_geometric.datasets. The Total Influence metric introduced in Section 4 is also released under torch_geometric.utils. We open-sourced the code for dataset construction and benchmarking at https://github.com/LeonResearch/City-Networks, and ensure that the experimental results and figures in this work are reproducible with the provided code.
Appendix A Network Statistics
@@
Since the exact calculation of $D$ has an impractical computational complexity at $\mathcal{O}(N^{2}\log(N))$, we use the following approach to estimate the approximate diameter $\hat{D}$ of our city networks. For all nodes in a given city, we select the ones with the maximum and minimum latitude and longitude, respectively: $coord(v_{1})=(\;\cdot\;,\;\text{lat}_{max})$, $coord(v_{2})=(\;\cdot\;,\;\text{lat}_{min})$, $coord(v_{3})=(\text{long}_{max},\;\cdot\;)$, and $coord(v_{4})=(\text{long}_{min},\;\cdot\;)$. Based on these, we compute the shortest path between $(v_{1},v_{2})$ and $(v_{3},v_{4})$, and take their maximum as our final diameter estimate, that is, $\hat{D}=\max\big(d(v_{1},v_{2}),d(v_{3},v_{4})\big)$. Note that the exact diameter $D$ will always be larger than our estimation $\hat{D}$.
@@
Appendix B Dataset Details
B.1 Raw Feature Processing
@@
The node and edge features in our dataset are derived from real-world features provided by OpenStreetMap for both road junctions and road segments, as detailed below.
@@
Feature engineering.
Since the categorical features land use, lanes, and road type contain many categories of only a few data points and varies across different networks, we only take the top $8$ categories with most entries for each categorical feature, and treat the rest as a single category - other. Based on our observations, we can cover more than 90% of the network with the top $8$ categories in all cases. This strategy leads to $12$ node features and $25$ edge features after one-hot encoding. Next, to facilitate a typical node classification task, we apply a simple neighborhood aggregation that transforms edge features into node features by averaging the features of incidental edges and then concatenating them to the features of the focal node. These new node features represent, for instance, the average speed limit around a road junction or the probability of finding a residential road nearby. As a result, the final dataset contains $37$ node features after processing. Lastly, we transform the graph into an undirected one by merging all edges between each pair of nodes into a single edge, where the edge features are averaged using to_undirected(reduce="mean") from PyG.
B.2 Long-range Node Labels based on Local Eccentricity
Controllable long-range signal based on local eccentricity.
As mentioned in Section 2.2, we use a $k$-hop ego-network to estimate the eccentricity of each node. Specifically, after obtaining the neighborhood $\mathcal{N}_{k}(v)$ within $k$ hops of node $v$, we compute the shortest path from node $v$ to all nodes within this neighborhood, and take the maximum as the local eccentricity $\hat{\varepsilon}_{k}(v)$:
$\hat{\varepsilon}_{k}(v)=\max_{u\in\mathcal{N}_{k}(v)}\rho_{w}(v,u),\quad\rho_{w}(v,u)=\min_{\pi\in P(v,u)}\sum_{e\in\pi}w(e),$ (12)
@@ B.2 Long-range Node Labels based on Local Eccentricity
where $\rho_{w}(v,u)$ is the weighted shortest path distance from node $v$ to node $u$, $P(v,u)$ denotes the set of all possible paths from node $v$ to node $u$, and $w(e)$ represents the edge weight for edge $e$. Here, we use one of the edge features road length as the edge weight $w(e)$, such that the local eccentricity $\hat{\varepsilon}_{k}(v)$ will indicate the maximal traveling distance (in meters) from node $v$ to its $k$-hop neighbors. Figure 5 shows the distributions of such approximation at $k=16$ across different city networks, in which Paris and Shanghai have the most skewed and uniform distributions, respectively.
Importantly, this method allows us to know a priori that the long-range signal should be highly correlated with the hop $k$, which facilitates us in task design and model benchmarking. Lastly, after obtaining the local eccentricity for all nodes, we split them into $10$ quantiles as the final node labels for transductive classification.
Computational cost.
[…]
Figure 6: Visualization of local eccentricity $\hat{\varepsilon}_{k}(v)$ at $k=[8,16,32]$ and full-hop (exact eccentricity) on two sub-regions. We can observe that in the last two cases (g) and (h), node labels become highly correlated with geographic coordinates and are hence less dependent on the graph structure.
B.3 The Roles of Graph Topology and Spatial Features
The importance of structural and spatial information.
Our goal is to use a label generation approach that incorporates long-range dependency via information from both node features and graph structures. Although the local eccentricity is based on a weighted shortest-path strategy with road length being the edge weight, it does not only capture a structure perspective. This is because road length naturally relates to geographical locations, road type, land use, etc., which will link the node labels to these node features. Since the edge features are removed for the transductive task, the model will not be able to directly infer the original weighted shortest-path values. Moreover, while we can control the range of the task by changing the approximation hop $k$, this knowledge is hidden from the model. Therefore, the model needs to explore the distant neighborhood and capture both structural and spatial information to effectively infer the long-range information.
Figure 7: An example graph in PascalVOC-SP from LRGB (Dwivedi et al., 2023), where nodes are represented by super-pixels and their labels are annotated by semantics. Although spatial coordinates are highly correlated with node labels, the inductive nature makes the task still graph-dependent.
Problem of uninformative graph structures on labels.
One catastrophic problem we encountered during the experiment is that, when using the exact eccentricity or a relatively larger $k$ for node labels, the signal becomes nearly independent of graph structures. As illustrated in Figure 6, we can observe a clear correlation between the node labels and their spatial coordinates, since nodes around the city center would generally have lower eccentricity values due to the 2D grid-like topology. This means that the spatial coordinates alone are sufficient for modelling $\varepsilon(v)$, hence weakening the importance of graph structure during modeling. Similarly, some other labeling approaches would also lead to the same problem, such as using a single node as an anchor, then regarding its shortest path distance to other nodes as the node label.
Finally, we would like to point out that this problem also exists in LRGB, as illustrated by an example graph from PascalVOC-SP in Figure 7. However, while the learning task in PascalVOC-SP resembles a spatial segmentation problem on the coordinates, it still requires the model to handle graph structural information for prediction due to the inductive task nature.
Figure 8: Results of GCN at different numbers of layers on Paris with $k=[8,16,32]$ used for labels.
B.4 Choice of the Long-range Level K
As explained in Section 2.2, the choice of $k$ is critical in designing the long-range signal. On one side, $k$ should be sufficiently large to distinguish our setting from short-range datasets (e.g., social graphs with typical message passing around $4$ hops). To further elaborate on this, we follow the same experiment setups in Section 3 and test GCN on Paris using $k=[8,16,32]$ at num_layers=$[2,4,8,16,32]$, where the results are reported in Figure 8 on the right.
We can observe that, at $k=8$, the model’s performance quickly saturates as expected when num_layers reaches $8$. While at $k=16$ and $32$, the performance starts to plateau at a much deeper depth of num_layers=$16$. As such, $8$ does not seem to be an optimal choice for $k$, since there is still room to “extend” the range of the signal.
On the other hand, $k$ should also remain small enough that it does not lead to the aforementioned problem of uninformative graph structure during modelling. As visualized in Figure 6, the underlying signal, given its local nature, becomes smoother as $k$ increases from $8$ to $32$. At the same time, we can also see a increasing correlation between signal and the spatial coordinates. In the extreme case when $k$ equals the network diameter, $\hat{\varepsilon}_{k}(v)$ becomes the exact eccentricity $\hat{\varepsilon}(v)$, and the classification task resembles a “segmentation” problem on a 2D plane. This also explains the marginal performance gain from num_layers=16 to 32 at $k=32$, as the signal becomes overly smoothed that information based on $16$-hop neighborhood is sufficient to have a good “guess” of the signal at hop $32$.
At the same time, for benchmarking purposes, we wish to test GNNs and GTs with enough model depth such that they can reach the source of the signal. However, we found that when training GTs on our large city networks, due to the quadratic complexity in their global attention mechanism, they often suffer from Out-Of-Memory errors (NVIDIA-L40@48GB ) when num_layers is large.
Therefore, we decided to adopt $k=16$ in our final dataset, which we believe achieves a balance between being sufficiently long-ranged and remaining practical for computation. As discussed earlier in Section 1 and Section 2.2, we believe City-Networks brings a new and important challenge to the literature, which calls for scalable architectures that can effectively handle long-range dependency.
Appendix C Experiment Details
In this section, we first provide training details in our experiments, and then present an ablation study that supports our main claim on the long-range dependency. Lastly, we summarize the best baseline results on our dataset and discuss potential limitations of our experimental settings.
C.1 Baselines and Datasets for Comparison
Baselines.
We consider four common GNNs: GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GCNII (Chen et al., 2020) for GNNs; and three Graph Transformers: GraphGPS (Rampášek et al., 2022), Exphormer (Shirzad et al., 2023), and SGFormer (Wu et al., 2023) as the baseline models in our experiments.
Datasets.
We compare the baseline results on City-Networks to the following datasets that are homophilic, heterophilic, large-scale, and long-range dependent.
  • Cora (Yang et al., 2016a) is a well-known homophilic citation network, with nodes representing documents and edges representing citation links. The features of the nodes are represented as bag-of-words that captures the content of the documents, and the goal is to predict the academic topic of each paper.
  • ogbn-arxiv (Hu et al., 2021) is a large-scale citation network of $169k$ computer science papers on arXiv that were indexed by the Microsoft academic graph, in which nodes represent papers and the directed edges indicate the citations. In particular, each node has an 128-dimensional feature vector derived from the word embeddings of titles and abstracts in the underlying papers. The task is to identify the primary category of each arXiv paper, that is, to classify each node into one of the 40 classes.
  • Amazon-Ratings (Platonov et al., 2023) is a heterophilic network that models Amazon product co-purchasing information, where nodes represent products and edges represent frequently co-purchased items. The goal is to predict average product ratings from five classes, with node features being the fastText embeddings of product descriptions.
  • PascalVOC-SP is an inductive dataset from LRGB (Dwivedi et al., 2023), which contains graphs derived from images. In particular, the nodes represent super-pixels and edges represent their boundaries. The labels are derived based on semantics, which makes the task similar to image segmentation.
As mentioned in Section 3, we closely follow the latest GNN tuning technique from Luo et al. (2024), which considers residual connection, batch normalization, dropout, etc., and use their code base for training classical GNNs and GTs on most datasets, except for PascalVOC-SP, where we adopt the hyperparameters reported by Tönshoff et al. (2024) and Shirzad et al. (2023).
C.2 Experimental Setups
As discussed earlier in Section 3, we consider transductive node classification with train/validation/test splits of $10\%/10\%/80\%$ on all four graphs in City-Networks.
Evaluation protocols. Since our goal is to investigate the presence of long-range dependency, we train each model at num_layers=[2, 4, 8, 16] while fixing hidden_size=128, and then check if the model’s performance will positively correlate with num_layers. All cases are repeated 5 times, and we present their means and standard deviations. In the ablation studies below, we also investigate different choices of hidden_size in [16, 32, 64, 128] following this strategy, which not only acts as hyperparameter tuning, but also shows the robustness of our conclusion to different choices of model hyperparameters.
Hyperparameters
We run each model for $30k$ maximum epochs using AdamW (Loshchilov & Hutter, 2019) optimizer for sufficient training, during which we record the validation accuracy every $100$ epochs and save the model at the best validation epoch for final testing. Meanwhile, we closely follow the latest GNN tuning technique from Luo et al. (2024), and our hyperparameter search space is summarized in Table 4. Note that for GTs, we do not apply positional encoding due to its impractical computation on our large graphs (which is one of the challenges of applying GTs on large graphs), and use GCN as their internal MPNNs.
Table 4: Summary of hyperparameters and their search space.
Type Hyper-parameter Search range Default
Model num_layers [2, 4, 8, 16][2,\ 4,\ 8,\ 16] $16$
hidden_size [16, 32, 64, 128][16,\ 32,\ 64,\ 128] $128$
pre_linear_layer [0, 1, 2][0,\ 1,\ 2] 0
post_linear_layer [0, 1, 2][0,\ 1,\ 2] 22
residual [True, False] True
Train learning rate $10^{-4},\ 5\times 10^{-4},\ 10^{-3}]$ $10^{-3}$
dropout [0, 0.2, 0.5, 0.7][0,\ 0.2,\ 0.5,\ 0.7] $0.2$
weight decay $0,\ 10^{-5},\ 5\times 10^{-5},\ 10^{-4}]$ $10^{-5}$
normalization [None, BatchNorm, LayerNorm] BatchNorm
C.3 Ablation Studies
Results under different hidden channel sizes.
To further support our main findings, we test two classical GNN baselines: GCN (Kipf & Welling, 2017) and GraphSAGE (Hamilton et al., 2017); and two GT baselines: Exphormer (Shirzad et al., 2023) and SGFormer (Wu et al., 2023) under different hidden_size=[16, 32, 64, 128] on Paris, Shanghai, and L.A.. The results are presented in Figure 9, where we can observe a consistent pattern with our main findings in Figure 3—the baselines’ performance improves substantially when increasing the number of layers from $2$ to $16$.
Notably, all baselines generally achieve superior performance with a larger hidden_size. However, such differences are often negligible compared to the increasing trend in model depths. With that being said, in the next section, we proceed to investigate scenarios when the model size is fixed.
Results at fix model depth with various sampling hops.
As discussed earlier, one limitation of the main results on Figure 3 is that the observed performance gains may be attributed to the increasing model parameters as the number of layers $L$ grows. To address this, we further investigate the scenario where $L$ is fixed at 16 (i.e., keeping the model size constant), and adopt an $H$-hop neighborhood sampling method introduced in GraphSAGE (Hamilton et al., 2017) that refrains the model from seeing beyond hop $H$. Note that the same method is also widely used in the literature for training models on large graphs.